A heavy block is suspended from a vertical spring.

The elastic potential energy is stored in the spring is 4 J.

What is the spring constant if the elongation of the spring is 10 cm?


1
Expert's answer
2021-06-11T11:22:15-0400

"\Delta l =10cm=0.1m"

"E_p=4\ J"

"E_p=\frac{k\Delta l^2}{2}"

"k= \frac{2E_p}{\Delta l^2}=\frac{2*4}{0.1^2}=0.08\ \frac{N}{m}"

"\text{Answer: }k=0.08\ \frac{N}{m}"